# Understanding Large Language Models in Your Pockets: Performance Study on COTS Mobile Devices

### v3

@@ Abstract
As large language models (LLMs) increasingly integrate into every aspect of our work and daily lives, there are growing concerns about user privacy, which push the trend toward local deployment of these models. There are a number of lightweight LLMs (e.g., Gemini Nano, LLAMA2 7B) that can run locally on smartphones, providing users with greater control over their personal data. As a rapidly emerging application, we are concerned about their performance on commercial-off-the-shelf mobile devices. To fully understand the current landscape of LLM deployment on mobile platforms, we conduct a comprehensive measurement study on mobile devices. While user experience is the primary concern for end-users, developers focus more on the underlying implementations. Therefore, we evaluate both user-centric metrics—such as token throughput, latency, and battery consumption—and developer-critical factors, including resource utilization, OS strategies, and inference engines. We also provide comprehensive comparisons across the mobile system-on-chips (SoCs) from major vendors, highlighting their performance differences in handling LLM workloads, which may help developers identify and address bottlenecks for mobile LLM applications. We hope that this study can provide insights for both the development of on-device LLMs and the design for future mobile system architecture.
Index Terms:
Large Language Models, Measurement, On device
@@ I Introduction
Over the past two years, Large Language Models (LLMs) have demonstrated exceptional performance across a wide range of domains. Their advanced capabilities in understanding human language and generating text have made them indispensable in various applications, including dialogue systems, machine translation, and text generation. Several companies have begun integrating LLMs as system services (e.g., Apple Intelligence [1], Copilot+ PC [2]), where these models assist users in organizing daily routines, summarizing emails and messages, and even generating automated replies. However, while LLMs serve as powerful and appealing personal assistants, these models continuously monitor and analyze users’ activities, which leads to serious privacy concerns [3].
@@
To address the privacy concerns, local deployment of LLMs is becoming attractive, as it allows these models to perform inference directly on the device and no personal data are sent to cloud servers. In this way, users have greater control of their personal data. While LLMs, such as ChatGPT [4] and Claude [5], are typically deployed on the cloud, local deployment are becoming feasible for two reasons. On one hand, a series of lightweight LLMs (e.g., Llama2-7B [6] and Mistral-7B [7]) have been introduced. These models, with billions of parameters, are significantly smaller in size compared to cloud-based models, which typically have hundreds of billions of parameters. Despite their reduced size, these lightweight LLMs maintain a high level of expressiveness in both common-sense reasoning and domain-specific knowledge, and thus can fully satisfy the basic needs of mobile users. On the other hand, state-of-the-art mobile SoCs demonstrate significant improvement in computing capability and memory capacity. More importantly, a variety of specialized AI accelerators are integrated into the mobile processors, such as Graphics Processing Units (GPUs) and Neural Processing Units (NPUs). These hardware improvements make it possible to implement complex AI functions on mobile devices.
@@
As a rapidly emerging application, LLMs differ significantly from traditional mobile applications in terms of computational complexity and memory demands. Therefore, we are particularly concerned with their performance on commercial off-the-shelf (COTS) mobile devices, as well as the performance fluctuations caused by mobile operating systems and hardware architecture compatibility. Several studies have taken the first steps to explore the current state of mobile LLM[8][9][10]. They have conducted preliminary assessments of model deployment and inference performance through the development of specialized benchmark suites. However, they mainly focus on performance metrics such as token throughput and latency, without looking into the underlying hardware dynamics and system configurations, such as CPU/GPU utilization, operating frequencies, etc. Besides, there is a lack of analysis about how these hardware and system-level factors affect on-device LLM inference. This is essential for developers to identify and address bottleneck for mobile LLM applications.
@@
In this paper, we aim to reveal the current state and potential of mobile hardware for supporting on-device LLMs by deploying LLMs across a diverse range of mobile devices. We select multiple high-tier SoCs from different vendors that represent currently available resources and development trends in mobile hardware. We analyze the differences in SoCs from major vendors and ultimately select SoCs represented by Qualcomm, HiSilicon, MediaTek and Apple to perform local LLM inference. We deploy a 77B model with $4$-bit quantization on mobile devices with llama.cpp [11] and MLC LLM [12], which are the two popular mobile LLM inference engines. During inference, we collect comprehensive metrics with specific profilers including Snapdragon Profiler [13], Arm Streamline [14] and Xcode [15] to make sure that all the data accurately reflect the hardware performance. We investigate how the hardware specifications, including big.LITTLE cores[16] and memory bandwidth, affect the inference performance, which is essential to help the development of future mobile SoCs to accelerate LLMs.
@@
Considering the diversity of mobile hardware and rapid iteration of capability, current research exhibits a notable gap in both comprehensive understanding and detailed evaluation of LLM deployment on mobile devices. In order to show the usability of on-device inference and provide developers with valuable insights, we have summarized the following key questions:
@@
Q2. Resource Configuration: What is the optimal runtime resource configuration, including system policies (DVFS and thermal throttling), the number of parallel threads, core affinity settings, and others? Moreover, how does the optimal configuration differ when inference runs exclusively versus concurrently with other tasks? Furthermore, how can this configuration be adapted to different devices?
Q3. Software-Hardware Alignment: Do the implementations and optimizations of inference frameworks align with general-purpose processors (CPUs) or specialized accelerators (GPUs) from various vendors? How much do software implementations limit the utilization of hardware capabilities, such as computation throughput or memory bandwidth?
@@
  • Performance of hardware accelerators: As for accelerators, Specialized AI accelerators, such as the Hexagon NPU, demonstrate remarkable prefill speed, achieving a $10$-fold increase compared to CPU and GPU on the same SoC. For decoding on GPU, due to the memory-bound nature of the decode stage, accelerating generation remains challenging.
  • RAM requirement: The memory (RAM) requirement to support LLMs depends on the model size and the quantization bit-width. A $4$-bit quantized 77B-parameter model typically requires at least $4$ GB of RAM to store the model weights and maintain a minimal context (eg. KV cache).
  • Battery support: For inference on modern high-tier mobile devices (e.g., Huawei Matepad12.612.6 Pro), the power consumption is around 88mAh per round. Typical smartphone batteries have a capacity of 40004000mAh to 60006000mAh, they can support hundreds of rounds of inference.
@@
  • Computing Thread: Given the “big.LITTLE” architecture of mobile CPUs, in most cases, setting the number of threads equal to the number of big cores (prime and performance cores) can achieve the optimal performance. Adding more cores (efficiency cores) may degrade the performance. When running in parallel with other tasks, the optimal configuration should vary dynamically, depending on both the real-time workload size and the nature of the co-running tasks.
Response to Q3:
  • Instruction speedup: Specialized machine instructions, such as smmla and sdot in Arm architecture, can significantly improve the performance. By carefully rearranging the weight matrix, these instructions can yield up to a 4× speed improvement.
@@ II-A Light-Weight LLMs
The number of LLMs has increased significantly in recent years. Numerous studies have explored the potential performance limits of these models by leveraging scaling laws [17] to increase their size. However, given the severe constraints of mobile platforms, it is crucial to compress these models to ensure that they are lightweight enough for deployment on edge devices. Current methods for compressing large models usually fall into the categories of pruning [18][19], distillation [20], quantization [21][22], and low-order decomposition [23]. Quantization is a widely used resource-efficient technique for compressing LLMs, reducing their memory and bandwidth requirements. The nature of quantization is a mapping from floating-point numbers to integers, especially the 4-bit integer quantization for mobile LLMs, which is recognized as the optimal compromise between model size and accuracy [9]. For example, based on the second-order information for weight updating, GPTQ [21] employs layer-wise quantization to minimize the increase in global loss. K-quant [24] is a group-quantization method that divides the weight matrix into multiple $16\times 8$ blocks, performing min-max quantization within each block. Due to outliers and dequantization overhead, most of the quantization algorithms focus on weight quantization only. However, by treating outliers carefully and implementing effective kernels, SmoothQuant [25] and OdesseyLLM [26] have tried weight-activation co-quantization to further lower the costs of LLM inference.
@@ II-B On-device Deployment of LLMs
Traditional machine learning (ML) and deep learning (DL) models have established a rich and mature ecosystem for mobile platforms. For example, TFLite, a DL inference engine designed for mobile devices, offers a wide range of highly optimized operators. Additionally, some vendors such as Qualcomm have developed hardware-specific DL libraries, which enhance the efficiency of model deployment and inference. However, unlike convolutional models, which typically have only a few hundred MB of parameters, LLMs feature orders of magnitude more parameters, more complex architectures, and higher computational demands. This complexity necessitates the development of specialized engines or operators specifically designed for efficient LLM inference.
There are some inference frameworks widely used including llama.cpp [11], MLC LLM [12], mnn-llm [32], mllm [33], etc. Among them,llama.cpp and mnn-llm are mainly optimized for LLM inference on CPUs, while MLC LLM leverages the computational capabilities of GPUs. Despite the availability of dedicated AI accelerators such as NPUs, APUs, and TPUs in the latest generation of mobile SoCs [34], developing for these accelerators presents challenges due to the typically closed-source nature of vendor-specific SDKs. Currently, only mllm and PowerInfer-2 [35] claim to support Qualcomm NPUs for accelerating LLM inference on mobile devices. The development of LLM solutions for AI accelerators with dedicated architectures is still an ongoing challenge for developers.
@@ II-C Existing Measurement Studies
However, none of these studies provide in-depth analysis about the impact of mobile hardware during inference, such as accelerator utilization, memory bandwidth, and the influence of DVFS on both CPU and GPU, heat buildup during multiple runs, parallel execution with other tasks. Although [9] touches on the impact of SoCs, it only examines the macro differences of SoCs from a single vendor. In contrast, our approach involves testing several SoCs from multiple vendors to gain a comprehensive understanding of how mainstream mobile processors support LLMs. This will help us identify hardware bottlenecks and potential future upgrades necessary for optimizing LLM performance on mobile devices.
@@ III-A Experimental Setup
Testing Model. We select Llama-2 77B as a representative model for mobile-compatible LLMs. Although most mobile devices have RAM capacities ranging from 88GB to 1616GB, the minimum effective available memory is around 44GB due to the substantial portion occupied by the operating system. Consequently, we determine that the 77B model size is the upper limit that most mobile hardware can support. Additionally, we employ the $4$-bit quantization scheme, which is widely used and provides an optimal balance between model size and accuracy.
Frameworks and Engines. In this paper, we would like to reveal the performance of LLM on a variety of different accelerators. To achieve this, we select two representative frameworks: llama.cpp for CPU deployments, and MLC LLM for general GPU deployments, to assess the inference capabilities of general accelerators. Additionally, llama.cpp supports GPU for specific SoCs, such as certain Snapdragon SoCs and Apple SoCs. For novel AI accelerators (e.g., NPU), due to the lack of available frameworks, rather than conducting our own device deployments, we refer to existing studies for results of mllm [38] and PowerInfer-$2$ [35], which are only available for Qualcomm NPUs. These references will be discussed in subsequent sections.
@@
  • MLC LLM [12] is a high-performance LLM deployment engine built on the TVM machine learning compiler, MLC LLM is designed to harness the power of mobile GPUs. Leveraging TVM, MLC LLM can automatically generate and optimize operators for a specific model and backend. These operators are then loaded and executed via the TVM Runtime. For Android-based GPUs, MLC LLM utilizes the OpenCL backend to ensure efficient execution.
@@
Evaluation Metrics. We collect results from two main aspects. First, we gather performance metrics that are of interest to users, including prefill and decoding speed, memory footprints, and battery consumption. These results are present in Section IV. Second, we analyze hardware dynamics, focusing on real-time processor utilization and working frequency, impact of Dynamic Voltage and Frequency Scaling (DVFS), heat accumulation and operator implementation, which developers are concerned about. These results are present in Section V. They provide clues for the performance metrics and can help us identify potential hardware and operating system bottlenecks during inference.
@@ III-B Workflow and Methodology
Measurement Methodology. To ensure accurate results, we perform two rounds of tests. In each round, a test is repeated five times, with an interval of $10$ seconds between each test. After the first round, the device reboots and rests for $10$ minutes to avoid overheating. This approach minimizes interference from other processes and reduces the impact of DVFS. The results are then averaged to provide a reliable measure of performance.

### v4

@@ Abstract
As large language models (LLMs) increasingly integrate into every aspect of our work and daily lives, there are growing concerns about user privacy, which push the trend toward local deployment of these models. There are a number of lightweight LLMs (e.g., Gemini Nano, LLAMA2 7B) that can run locally on smartphones, providing users with greater control over their personal data. As a rapidly emerging application, we are concerned about their performance on commercial-off-the-shelf mobile devices. To fully understand the current landscape of LLM deployment on mobile platforms, we conduct a comprehensive measurement study on mobile devices.
While user experience is the primary concern for end-users, developers focus more on the underlying implementations. Therefore, we evaluate both user-centric metrics—such as token throughput, latency, and response quality—and developer-critical factors, including resource utilization, OS strategies, battery consumption, and launch time. We also provide comprehensive comparisons across the mobile system-on-chips (SoCs) from major vendors, highlighting their performance differences in handling LLM workloads, which may help developers identify and address bottlenecks for mobile LLM applications. We hope that this study can provide insights for both the development of on-device LLMs and the design for future mobile system architecture.
@@ I Introduction
Over the past two years, Large Language Models (LLMs) have demonstrated exceptional performance across a wide range of domains. Their advanced capabilities in understanding human language and generating text have made them indispensable in various applications, including dialogue systems, machine translation, and text generation. Several companies have begun integrating LLMs as system services (e.g., Apple Intelligence [29], Copilot+ PC [43], where these models assist users in organizing daily routines, summarizing emails and messages, and even generating automated replies. However, while LLMs serve as powerful and appealing personal assistants, these models continuously monitor and analyze users’ activities, which leads to serious privacy concerns [26].
@@
To address the privacy concerns, local deployment of LLMs is becoming attractive, as it allows these models to perform inference directly on the device and no personal data are sent to cloud servers. In this way, users have greater control of their personal data. While LLMs, such as ChatGPT [1] and Claude [3], are typically deployed on the cloud, local deployment is becoming feasible for two reasons. On one hand, with the development of resource-efficient strategies [58], a series of lightweight LLMs (e.g., Llama2-7B [54] and Mistral-7B [30]) have been introduced. These models, with billions of parameters, are significantly smaller in size compared to cloud-based models, which typically have hundreds of billions of parameters. Despite their reduced size, these lightweight LLMs maintain a high level of expressiveness in both common-sense reasoning and domain-specific knowledge, and thus can fully satisfy the basic needs of mobile users. On the other hand, state-of-the-art mobile SoCs demonstrate significant improvement in computing capability and memory capacity. More importantly, a variety of specialized AI accelerators are integrated into the mobile processors, such as Graphics Processing Units (GPUs) and Neural Processing Units (NPUs). These hardware improvements make it possible to implement complex AI functions on mobile devices.
@@
As a rapidly emerging application, LLMs differ significantly from traditional mobile applications in terms of computational complexity and memory demands [63]. Therefore, we are particularly concerned with their performance on commercial off-the-shelf (COTS) mobile devices, as well as the performance fluctuations caused by mobile operating systems and hardware architecture compatibility. Several studies have taken the first steps to explore the current state of mobile LLM[33][35][46]. They have conducted preliminary assessments of model deployment and inference performance through the development of specialized benchmark suites. However, they mainly focus on performance metrics such as token throughput and latency, without looking into the underlying hardware dynamics and system configurations, such as CPU/GPU utilization, operating frequencies, etc. Besides, there is a lack of analysis about how these hardware and system-level factors affect on-device LLM inference. This is essential for developers to identify and address bottlenecks for mobile LLM applications.
@@
In this paper, we aim to reveal the current state and potential of mobile hardware for supporting on-device LLMs by deploying LLMs across a diverse range of mobile devices. We select multiple high-tier SoCs from different vendors that represent currently available resources and development trends in mobile hardware. We analyze the differences in SoCs from major vendors and ultimately select SoCs represented by Qualcomm, HiSilicon, MediaTek and Apple to perform local LLM inference. We deploy six models from various families (Llama22/$3.2$ [44], Qwen33 [60], Gemma33 [51]) on mobile devices with llama.cpp [20] and MLC LLM [52], which are the two popular mobile LLM inference engines. During inference, we collect comprehensive metrics with specific profilers including Snapdragon Profiler [47], Arm Streamline [6] and Xcode [5] to make sure that all the data accurately reflect the hardware performance. We investigate how the hardware specifications, including big.LITTLE cores[14] and accelerators’ architecture, affect the inference performance, which is essential to help the development of future mobile SoCs to accelerate LLMs.
@@
Considering the diversity of mobile hardware and rapid iteration of capabilities, current research exhibits a notable gap in both comprehensive understanding and detailed evaluation of LLM deployment on mobile devices. In order to show the performance of on-device inference and provide developers with valuable insights, we have summarized the following key questions:
@@
Q2. Model and Quantization: What is the trade-off between model accuracy and throughput across different quantization schemes? How does model architecture influence sensitivity to quantization?
Q3. Resource Configuration: What is the optimal runtime resource configuration, including system policies (DVFS and thermal throttling), the number of parallel threads, core affinity settings, and others? Moreover, how does the optimal configuration differ when inference runs exclusively versus concurrently with other tasks? Furthermore, how can this configuration be adapted to different devices?
Q4. Software-Hardware Alignment: Do the implementations and optimizations of inference frameworks align with general-purpose processors (CPUs) or specialized accelerators (GPUs) from various vendors? How much do software implementations limit the utilization of hardware capabilities, such as computation throughput or memory bandwidth?
@@
  • Performance of hardware accelerators: As for accelerators, GPU demonstrates promising preliminary acceleration in the prefill stage, while this speedup is limited in the memory-bound decode stage. At this stage, GPU acceleration support is available only for a limited set of devices.
  • RAM requirement: The memory (RAM) requirement to support LLMs depends on the model size and the quantization bit-width. For most mobile devices, a $4$-bit quantized 77B-parameter model is the upper limit. It typically requires at least $4$ GB of RAM to store the model weights and maintain a minimal context (e.g., KV cache).
  • Battery support: For CPU-based inference with a $64$-token prompt and $128$-token generation on modern high-tier mobile devices, power consumption is approximately $20$ mAh per round. Given that typical smartphone batteries have a capacity of $4000$ mAh to $6000$ mAh, these devices can support hundreds of rounds of inference. Instruction acceleration for CPU inference further improves power efficiency. Notably, GPU-based inference provides a significant energy efficiency advantage over CPU execution, highlighting its potential for optimized performance on mobile devices.
@@
  • Trade-off between accuracy and throughput: 4-bit quantization strikes the optimal balance by achieving peak or near-peak speeds in both prefill and decoding stages while maintaining robust inference accuracy. For small-scale LLMs, such as Gemma33-11B, 88-bit quantization serves as a viable alternative to enhance response quality without sacrificing throughput.
  • Model Architecture and Sensitivity to Quantization: Most models maintain performance comparable to the F16 baseline across various bit-widths, with significant degradation occurring only at extremely low levels, such as 3-bit. However, there is an exception—higher bit-widths do not necessarily improve accuracy for some models on certain tasks. For example, on TruthfulQA [38], Llama22-77B and Gemma33-11B experience accuracy declines when transitioning from Q33_K_M to Q44_K_M.
Response to Q3:
@@
  • Computing Thread: Given the “big.LITTLE” architecture of mobile CPUs, in most cases, setting the number of threads equal to the number of big cores (prime and performance cores) can achieve the optimal performance. Adding more cores (efficiency cores) may degrade the performance. When running in parallel with other tasks, the optimal configuration should vary dynamically, depending on both the real-time workload and the nature of the co-running tasks.
Response to Q4:
  • Instruction speedup: Specialized machine instructions, such as smmla and sdot in Arm architecture, can significantly improve the performance. By carefully rearranging the weight matrix, these instructions can yield up to a 6× speed improvement for Llama22-77B.
@@ II-A Light-Weight LLMs
The number of LLMs has increased significantly in recent years. Numerous studies have explored the potential performance limits of these models by leveraging scaling laws [31] to increase their size. However, given the severe constraints of mobile platforms and the high deployment costs [57] [36], it is crucial to compress these models to ensure that they are lightweight enough for deployment on edge devices. Current methods for compressing large models usually fall into the two principal approaches: algorithmic-level and infrastructure-level [36]. The algorithmic-level optimization includes pruning [40][55], distillation [53], low-order decomposition [59], and efficient attention implementation [62] [17]. The infrastructure-level optimizations are exemplified by quantization [19][37]. Quantization is a widely used resource-efficient technique for compressing LLMs, reducing their memory and bandwidth requirements. The nature of quantization is a mapping from floating-point numbers to integers, especially the 4-bit integer quantization for mobile LLMs, which is recognized as the optimal compromise between model size and accuracy [35]. For example, based on the second-order information for weight updating, GPTQ [19] employs layer-wise quantization to minimize the increase in global loss. K-quant [21] is a group-quantization method that divides the weight matrix into multiple $16\times 8$ blocks, performing min-max quantization within each block. Due to outliers and dequantization overhead, most of the quantization algorithms focus on weight quantization only. However, by treating outliers carefully and implementing effective kernels, SmoothQuant [56] and OdesseyLLM [34] have tried weight-activation co-quantization to further lower the costs of LLM inference.
@@ II-B On-device Deployment of LLMs
Traditional machine learning (ML) and deep learning (DL) models have established a rich and mature ecosystem for mobile platforms. For example, TFLite, a DL inference engine designed for mobile devices, offers a wide range of highly optimized operators. Additionally, some vendors such as Qualcomm have developed hardware-specific DL libraries, which enhance the efficiency of model deployment and inference. However, unlike convolutional models, which typically have only a few hundred MB of parameters, LLMs feature orders of magnitude more parameters, more complex architectures, and higher computational demands. Although recent studies [39] have demonstrated the efficient deployment of Transformer-based models at the edge, models with only millions of parameters exhibit a significant gap compared to contemporary billion-parameter LLMs in terms of both performance and deployment complexity. This complexity requires the development of specialized engines or operators to reduce latency and memory management strategies to minimize memory costs, all specifically designed for efficient LLM inference.
There are some inference frameworks widely used including llama.cpp [20], MLC LLM [52], mnn-llm [2], mllm [61], etc. Among them, llama.cpp and mnn-llm are mainly optimized for LLM inference on CPUs, while MLC LLM leverages the computational capabilities of GPUs. Despite the availability of dedicated AI accelerators such as NPUs, APUs, and TPUs in the latest generation of mobile SoCs [28], developing on these accelerators presents challenges due to the closed-source nature of vendor-specific SDKs.
@@ II-C Existing Measurement Studies
However, none of these studies provide in-depth analysis about the impact of mobile hardware during inference, such as accelerator utilization, memory bandwidth, and the influence of DVFS on both CPU and GPU, heat buildup during multiple runs, parallel execution with other tasks. Although [35] touches on the impact of SoCs, it only examines the macro differences of SoCs from a single vendor. In contrast, this study involves testing several SoCs from multiple vendors to gain a comprehensive understanding of how mainstream mobile processors support LLMs. This will help us identify hardware bottlenecks and potential future upgrades necessary for optimizing LLM performance on mobile devices.
@@ III-A Experimental Setup
Testing Model. We evaluate six representative LLMs from the Llama, Qwen, and Gemma families, with parameter sizes ranging from 11B to 77B (Table II). This selection represents a range of mobile application scenarios, from lightweight models for tasks like message summarization to larger models designed for autonomous agents. Since most mobile devices come with 88GB to 1616GB of RAM, with at least 44GB available after accounting for the operating system, 77B-model is the upper limit for hardware compatibility. By default, we employ Llama22-77B with $4$-bit quantization as the evalutaion model.
Frameworks and Engines. In this paper, we would like to reveal the performance of LLM on a variety of different accelerators. To achieve this, we select two representative frameworks: llama.cpp for CPU deployments, and MLC LLM for general GPU deployments, to assess the inference capabilities of general accelerators. Additionally, llama.cpp supports GPU for specific SoCs, such as certain Snapdragon SoCs and Apple SoCs.
@@
  • MLC LLM [52] is a high-performance LLM deployment engine. Built on the TVM machine learning compiler, MLC LLM is designed to harness the power of mobile GPUs. Leveraging TVM, MLC LLM can automatically generate and optimize operators for a specific model and backend. These operators are then loaded and executed via the TVM Runtime. For Android-based GPUs, MLC LLM utilizes the OpenCL backend to ensure efficient execution.
@@ III-B Workflow and Methodology
Measurement Methodology. For CPU-based inference, in order to minimize the effects of DVFS and ensure reliable results, we conduct two rounds of testing. In every round, each test (one round of inference) is repeated five times, with a 10-second interval between repetitions. After the first round, the device is rebooted and allowed to rest for 10 minutes to prevent overheating. For GPU-based inference, since the GPU frequency is typically stable, we perform five tests with a minimum 10-minute rest period between each test. The results are then averaged to provide a reliable measure of performance.
TABLE II: LLM model specifications
Model Name Model Size Detailed Version (Huggingface)
Llama3.23.2-11B 1.231.23B meta-llama/Llama-3.23.2-11B-Instruct
Llama3.23.2-33B 3.213.21B meta-llama/Llama-3.23.2-33B-Instruct
Llama22-77B 77B meta-llama/Llama-22-77b-chat-hf
Qwen33-1.71.7B 1.71.7B Qwen/Qwen33-1.71.7B
Qwen33-44B 4.04.0B Qwen/Qwen33-44B-Instruct-$2507$
Gemma33-11B 1.01.0B google/gemma-33-11b-it